Find all the asymptotes of the curve,

x³+y³ = 3axy, where a>0.

please solve it very urgent
1
Expert's answer
2019-10-30T04:43:55-0400

Let's find a slant asymptote of this curve

Let's rewrite the equation

"x^3+y^3-3axy=0"

The slant asymptote is "y=mx+c"

Substuting this to the equation will give us

"x^3+(mx+c)^3-3ax(mx+c)=(1+m^2)x^3+(3m^2c+3am)x^2+(3mc^2-3ac)x+c^3=0"

Therefore

"1+m^3=0;\qquad m=-1"

"3m^2c-3am=0;\qquad mc-a=0;\qquad c=-a"

Hence "y=-x-a" is a slant asymptote of this curve